Question 15 You are valuing the data processing company KirsebergFact whose Free Cash-Flows are projected to evolve in accordance with the following table: Year 2 30 FCF (Millions SEK) 16 32 30 The weighted Average Cost of Capital is equal to 10% and the FCFs after year four are forecasted to grow at the industry average of 3.5%. Furthermore, the company has 200 million SEK in debt and excess cash equal to 5 million SEK. If its number of outstanding shares is equal to 35 million, what do you estimate the share price of the company to be closest to A. 6.15 B. 5.25

hi guys i know how to solve this i just dont understand why in year 4 you discount year 4 cash flow and year 4 price with 1,1^3 when its year 4?
Answer #1

FCF1 = 16
FCF2 = 30
FCF3 = 32
FCF4 = 30

Growth rate, g = 3.50%
WACC = 10%

FCF5 = FCF4 * (1 + g)
FCF5 = 30 * 1.035
FCF5 = 31.05

Horizon Value, V4 = FCF5 / (WACC - g)
Horizon Value, V4 = 31.05 / (0.10 - 0.035)
Horizon Value, V4 = 31.05 / 0.065
Horizon Value, V4 = 477.6923

Value of Firm, V0 = 16/1.10 + 30/1.10^2 + 32/1.10^3 + 30/1.10^4 + 477.6923/1.10^4
Value of Firm, V0 = 410.14

Value of Firm = 410.14 million SEK

Value of Equity = Value of Firm - Value of Debt + Cash
Value of Equity = 410.14 million SEK - 200 million SEK + 5 million SEK
Value of Equity = 215.14 million SEK

Share Price = Value of Equity / Number of shares
Share Price = 215.14 million SEK / 35 million
Share Price = 6.15 SEK